Tyler has 4 times as many coins as Oliver. If Oliver had 27 more coins, the two boys would have exactly the same amount. How many coins does Tyler have?
Answer by ikleyn(52795)

.
Tyler has 4 times as many coins as Oliver. If Oliver had 27 more coins,
the two boys would have exactly the same amount. How many coins does Tyler have?
~~~~~~~~~~~~~~~~~~

 x = # of coins Oliver has;

4x = # of coins Tyles has.


Equation

    x + 27 = 4x.


Solution

    27 = 4x - x

    27 = 3x

     x = 27/3 = 9.


ANSWER.  Tyler has 4*9 = 36 coins.

Solved.